# A geometric phase approach to quark confinement from stochastic gauge-geometry flows

### v1

@@ Appendix A $3$-Manifold topology and geometry
We hereby collect some general facts on $3$-manifolds and their geometry. A connected 3-manifold $N$ is prime if it cannot be obtained as a connected sum of two manifolds $N_{1}\#N_{2}$ neither of which is the 3-sphere $S^{3}$ (or, equivalently, neither of which is the homeomorphic to $N$). Examples are the 3-torus $T^{3}$ and $S^{1}\times S^{2}$ but also the Poincaré sphere. According to Milnor (1962), any compact, oriented 3-manifold is the connected sum of a unique (up to homeomorphism) collection of prime 3-manifolds (prime decomposition). A subset of prime manifolds are the irreducible 3-manifolds. A connected 3-manifold is irreducible if every differentiable submanifold $S$ homeomorphic to a sphere $S^{2}$ bounds a subset $D$ (i.e. $\partial D=S$) which is homeomorphic to the closed ball $D^{3}$. Moreover, it is known that the only (orientable) prime but reducible 3-manifold is $S^{1}\times S^{2}$.
To understand the geometric properties, a finer decomposition induced by incompressible tori is needed. A properly embedded connected surface $S\subset N$ is called 2-sided if its normal bundle is trivial, and 1-sided if its normal bundle is nontrivial. A 2-sided connected surface $S$ other than $S^{2}$ or $D^{2}$ is called incompressible if for each disk $D\subset N$ with $D\cap S=\partial D$ there is a disk $D^{\prime}\subset S$ with $\partial D^{\prime}=\partial D$. The boundary of a 3-manifold is an incompressible surface. Most importantly, the 3-sphere $S^{3}$, $S^{2}\times S^{1}$ and the 3-manifolds $S^{3}/\Gamma$ with $\Gamma\subset SO(4)$ a finite subgroup do not contain incompressible surfaces. The class of 3-manifolds $S^{3}/\Gamma$ (the spherical 3-manifolds) include cases like the Poincaré sphere ($\Gamma=I^{*}$ the binary icosahedral group) or lens spaces ($\Gamma=\mathbb{Z}_{p}$ the cyclic group).
@@
where #$\#_{n}$ denotes the $n$-fold connected sum and $\Gamma\subset SO(4)$ is a finite subgroup. The decomposition of $N$ is unique up to the order of the factors. The irreducible 3-manifolds $K_{1},\ldots,\,K_{n_{1}}$ are able to contain incompressible tori and one can split $K_{i}$ along the tori into simpler pieces $K=H\cup_{T^{2}}G$ Jaco and Shalen (1979) (called the JSJ decomposition). The two classes $G$ and $H$ are the graph manifold $G$ and the hyperbolic 3-manifold $H$.
The hyperbolic 3-manifold $H$ has a torus boundary $T^{2}=\partial H$, i.e. $H$ admits a hyperbolic structure in the interior only. It is also useful to consider the splitting of the link/knot complement. As shown in Budney (2005), the Whitehead double of a knot leads to JSJ decomposition of the complement into the knot complement and the complement of the Whitehead link (along one torus boundary of the Whitehead link complement).
One property of hyperbolic 3-manifolds is central for our considerations: the Mostow rigidity. As shown by Mostow Mostow (1968), every hyperbolic $n-$manifold $n>2$ with finite volume has the property that Every diffeomorphism (especially, every conformal transformation) is induced by an isometry. Therefore, one cannot scale a hyperbolic 3-manifold and the volume is a topological invariant. Together with the prime and JSJ decomposition,
@@
we can discuss the geometric properties central to Thurstons geometrization theorem: every oriented closed prime 3-manifold can be cut along tori (JSJ decomposition), so that the interior of each of the resulting manifolds has a geometric structure with finite volume. Now, we have to clarify the term ‘geometric structure’. A model geometry is a simply connected smooth manifold $X$ together with a transitive action of a Lie group $G$ on $X$ with compact stabilizers.

### v2

@@ Appendix A $3$-Manifold topology and geometry
We hereby collect some general facts on $3$-manifolds and their geometry. A connected 3-manifold $N$ is prime if it cannot be obtained as a connected sum of two manifolds $N_{1}\#N_{2}$, neither of which is the 3-sphere $S^{3}$ — or, equivalently, neither of which is homeomorphic to $N$. Examples are the 3-torus $T^{3}$ and $S^{1}\times S^{2}$, as well as the Poincaré sphere. According to Milnor (1962), any compact, oriented 3-manifold is the connected sum of a unique (up to homeomorphism) collection of prime 3-manifolds (prime decomposition). A subset of prime manifolds are the irreducible 3-manifolds. A connected 3-manifold is irreducible if every differentiable submanifold $S$ homeomorphic to a sphere $S^{2}$ bounds a subset $D$ (i.e. $\partial D=S$) that is homeomorphic to the closed ball $D^{3}$. Moreover, it is known that the only (orientable) prime but reducible 3-manifold is $S^{1}\times S^{2}$.
To understand the geometric properties, a finer decomposition induced by incompressible tori is needed. A properly embedded connected surface $S\subset N$ is called 2-sided if its normal bundle is trivial, and 1-sided if its normal bundle is nontrivial. A 2-sided connected surface $S$ other than $S^{2}$ or $D^{2}$ is called incompressible if for each disk $D\subset N$ with $D\cap S=\partial D$ there is a disk $D^{\prime}\subset S$ with $\partial D^{\prime}=\partial D$. The boundary of a 3-manifold is an incompressible surface. Most importantly, the 3-sphere $S^{3}$, $S^{2}\times S^{1}$ and the 3-manifolds $S^{3}/\Gamma$ with $\Gamma\subset SO(4)$ a finite subgroup do not contain incompressible surfaces. The class of 3-manifolds $S^{3}/\Gamma$ (the spherical 3-manifolds) include cases like either the Poincaré sphere ($\Gamma=I^{*}$ the binary icosahedral group) or lens spaces ($\Gamma=\mathbb{Z}_{p}$ the cyclic group).
@@
where #$\#_{n}$ denotes the $n$-fold connected sum and $\Gamma\subset SO(4)$ is a finite subgroup. The decomposition of $N$ is unique up to the order of the factors. The irreducible 3-manifolds $K_{1},\ldots,\,K_{n_{1}}$ are able to contain incompressible tori. So one can split $K_{i}$ along the tori into simpler pieces $K=H\cup_{T^{2}}G$ Jaco and Shalen (1979) (called the JSJ decomposition). The two classes $G$ and $H$ are the graph manifold $G$ and the hyperbolic 3-manifold $H$.
The hyperbolic 3-manifold $H$ has a torus boundary $T^{2}=\partial H$, i.e. $H$ admits a hyperbolic structure in the interior only. It is also useful to consider the splitting of the link/knot complement. As shown in Budney (2005), the Whitehead double of a knot leads to the JSJ decomposition of the complement into the knot complement and the complement of the Whitehead link (along one torus boundary of the Whitehead link complement).
One property of hyperbolic 3-manifolds is central for our considerations: the Mostow rigidity. As shown by Mostow Mostow (1968), every hyperbolic $n-$manifold $n>2$ with finite volume has the property that every diffeomorphism (especially, every conformal transformation) is induced by an isometry. Therefore, one cannot scale a hyperbolic 3-manifold, and thus the volume is a topological invariant. Together with the prime and the JSJ decomposition,
@@
we may then discuss the geometric properties central to he Thurstons geometrisation theorem: every oriented closed prime 3-manifold can be cut along tori (JSJ decomposition), so that the interior of each of the resulting manifolds has a geometric structure with finite volume. We hereby clarify the term ‘geometric structure’. A model geometry is a simply connected smooth manifold $X$ together with a transitive action of a Lie group $G$ on $X$ with compact stabilisers.
@@
Appendix B Embedding Casson handle by using quadratic differentials and vertical foliations
A rigorous mathematical argumentation is based on the embedding of the tree $T_{CH}$ for a Casson handle $CH$. The Casson handle is a branched surface and it can be described by quadratic differentials as follows. Let $X$ be a Riemannian surface then a quadratic differential is a section of $T^{*}X^{1,0}\otimes T^{*}X^{1,0}$ which is locally given by
$q=q(z)dz^{2}=q(z)\,dz\otimes dz$
with the holomorphic function $q(z)$. Away from the zeros of $q(z)$ we can choose a canonical conformal coordinate $\xi(z)=\intop^{z}\sqrt{q}$ so that $q=d\xi^{2}$. Then the set $\left\{Re(\xi)={\rm const.}\right\}$ defines a foliation, called the vertical measured foliation. The holomorphic function $q(z)$ can be locally expressed as a polynomial. The zeros of the polynomial are the branching points of the surface, i.e. $z^{n}$ branches into $n$ pieces at $z=0$. Using this result, we are able to generate the tree $T_{CH}$ by a polynomial. Furthermore we identify the coordinate $t$ with $t=Re(\xi)$ so that $t=const.$ defines a vertical foliation (into slices of the constant time). By the theorem of Hubbard and Masur Hubbard and Masur (1979), for every measured foliation on $X$ (of genus $g>1$) there exists a unique quadratic differential so that its vertical measured foliation is equivalent to the measured foliation. In our case, the infinite tree seen as branched surface is the covering space of a Riemannian surface of infinite genus. Therefore the quadratic differential is unique by lifting it to the covering.